# Randomized Controlled Trials for Security Copilot for IT Administrators

### v1

@@ Abstract
As generative AI (GAI) tools become increasingly integrated into workplace environments, it is essential to measure their impact on productivity across specific domains. This study evaluates the effects of Microsoft’s Security Copilot (“Copilot”) on information technology administrators (“IT admins”) through randomized controlled trials. Participants were divided into treatment and control groups, with the former granted access to Copilot within Microsoft’s Entra and Intune admin centers. Across three IT admin scenarios – sign-in troubleshooting, device policy management, and device troubleshooting – Copilot users demonstrated substantial improvements in both accuracy and speed. Across all scenarios and tasks, Copilot subjects experienced a 34.53% improvement in overall accuracy and a 30.69% reduction in task completion time. We also find that the productivity benefits vary by task type, with more complex tasks showing greater improvement. In free response tasks, Copilot users identified 147.07% more relevant facts and reduced task completion time by 61.94%. Subject satisfaction with Copilot was high, with participants reporting reduced effort and a strong preference for using the tool in future tasks. These findings suggest that GAI tools like Copilot can significantly enhance the productivity and efficiency of IT admins, especially in scenarios requiring information synthesis and complex decision-making.
@@ 1 Introduction
As generative AI (GAI) tools continue to mature, organizations are interested in understanding their impact on workforce productivity to guide adoption decisions. However, the effects of these tools can vary significantly depending on the specific domain and application. So, it is important to study each application individually. In this study, we evaluate the benefits of generative AI on information technology administrators (“IT admins”) by conducting a randomized controlled trial (RCT) with IT admins using Microsoft’s Security Copilot (“Copilot”). We find that Copilot confers substantial benefits to IT admins, including accuracy improvements across all scenarios and tasks (34.53%*** on average) and time savings across all scenarios (30.69%*** on average). [^1] Throughout, we indicate statistical significance using asterisks, where * indicates $0.05<p\leq 0.1$, ** indicates $0.01<p\leq 0.05$, and *** indicates $0<p\leq 0.01$.
@@
We recruited 182 subjects and asked them to sign into an environment created for this experiment. Half of the subjects were assigned to the control group and were granted access to the Entra and Intune admin centers. Half of subjects were assigned to the treatment group and were additionally given access to Copilot within the Entra and Intune admin centers. Then, they used these tools and the simulated data in that environment to respond to our experimental tasks.
@@ 3 Methodology
We recruited our subject pool through Upwork, a marketplace for freelancers. We targeted freelancers with a mix IT admin experience. We required subjects to be proficient in reading and writing English and to have a positive reputation on Upwork. We told them to expect the tasks to take approximately two hours to complete. We offered performance incentives for combined speed and accuracy. Specifically, we multiplied the percentiles of each subject’s speed and accuracy and awarded payments on the following scale:
@@ 4 Results
The results answer RQ1 in the affirmative: Copilot does offer speed and accuracy improvements in common IT scenarios. Across all tasks, we measure a 34.53%*** increase in accuracy and a 30.69%*** reduction in task completion time. In the subsections on accuracy and speed below, we present detailed results answering RQ1-RQ4.
@@ 4.1 Accuracy
Free Response Content 3 1.66 0.67 147.07%***
@@
Copilot shows statistically significant accuracy improvements in two out of three task types. The effect sizes generally track the task complexity, with the greatest benefits arising in free response tasks where Copilot subjects reported 147.07%*** more relevant facts than our control group. In contrast to MC and SATA tasks, which only require selecting the correct answers from a short list of possibilities, the free response tasks leverage AI’s unique ability to find and report relevant facts about a given scenario from a large and complex corpus, e.g., looking through sign-in logs to find and categorize sign-in failures caused by conditional access policies. Overall, Copilot yields an accuracy improvement of 21.76%*** on SATA tasks across the sign-in troubleshooting and device troubleshooting scenarios.
@@ 4.2 Speed
We measure statistically significant time savings in all scenarios. Table 3 summarizes these results by task type and scenario. The “Aggregate Results, All Content (Holding accuracy constant)” numbers represent our headline speed improvements, a 29.10%*** time savings for Copilot users. We also note that Copilot generally involves greater latency than typical page-loading in the admin centers, which is common among GAI tools. This necessarily slowed the Copilot users. Product improvements should reduce this latency and further increase the time savings for users with Copilot.
@@
All Content 53.74 44.29 62.52 -29.10%***
SATA 17.77 15.25 20.12 -24.20%***
MC 10.74 12.02 9.55 25.78%***
MC (Points $\geq$ 1) 10.80 12.28 9.54 28.75%***
Free Response 31.38 17.03 44.73 -61.94%***
@@
All Content 24.00 16.63 30.86 -46.11%***
SATA 6.38 5.04 7.64 -34.07%***
Free Response Content 14.85 11.59 22.57 -48.62%***
@@
All Content 19.90 17.45 22.17 -21.31%***
MC 10.74 12.02 9.55 25.78%***
MC (Points $\geq$ 1) 10.80 12.28 9.54 28.75%***
Free Response Content 12.65 5.43 19.36 -71.95%***
@@
We measure time savings for Copilot across all three scenarios. The time savings included 46.11%*** for sign-in troubleshooting, 21.31%*** for device policy management, and 15.69% *** for device troubleshooting.
@@
We find that Copilot’s time savings track its accuracy improvements across task types. That is, the biggest gains again come from free response (61.94%*** time savings), followed by SATA (24.20%***). Lastly, the MC questions show a decrease in speed for Copilot users (25%*** time increase). For these MC questions we also estimate a negative relationship between accuracy and time spent for the control group. This is likely because control subjects who knew the answer responded quickly, and those who did not know it gained little accuracy by spending more time by, for example, searching the internet. This can explain the data we see, where even though Copilot provided an advantage in responding accurately, controlling for accuracy in calculating its time savings does not make up for the disadvantage of its latency.
We also notice that Copilot users spent much more time than the control users on the first MC question compared to the subsequent three questions (see figure 2). We think this suggests Copilot users spent extra time on the first question “learning” how to use Copilot for this type of task. Therefore, we drop the first question and recompute the time differences across the remaining tasks. The result is a milder 15.44%* increase in task completion time for Copilot users.
@@
As we did with accuracy, we examine time savings across different experience levels and domains. Our findings once again reveal statistically significant point estimates of time savings for all groups, as shown in figure 3. However, we do not observe statistically significant differences in estimated effect sizes between any of the groups. The left panel illustrates Copilot’s impact on task completion times for the sign-in troubleshooting task, with a 51.5%*** and 44.4%*** time savings for Copilot users with 0-3 and 3+ years of experience, respectively. These point estimates are consistent with intuition and our accuracy findings.
The right panel of figure 3 shows time savings of 12.0%*** and 29.7%*** for 0-3 and 3+ years of experience, respectively, for device management tasks. This suggests that experienced network management subjects benefit more from Copilot in terms of time savings than their less experienced counterparts, which runs counter to intuition and expectation. However, there is an explanation. This is driven by the fact that less experienced control subjects exhibit a negative relationship between task completion time and accuracy for the device policy management MC question, which shrinks the possible time savings for the less experienced Copilot subjects. The reason for the negative relationship is likely that less experienced control users fell into two categories: those that knew the answer quickly and those that did not know the answer and went looking. Those that went looking were largely unsuccessful in finding the right answer for this set of MC questions.

### v2

@@ Abstract
As generative AI (GAI) tools become increasingly integrated into workplace environments, it is essential to measure their impact on productivity across specific domains. This study evaluates the effects of Microsoft’s Security Copilot (“Copilot”) on information technology administrators (“IT admins”) through randomized controlled trials. Participants were divided into treatment and control groups, with the former granted access to Copilot within Microsoft’s Entra and Intune admin centers. Across three IT admin scenarios – sign-in troubleshooting, device policy management, and device troubleshooting – Copilot users demonstrated substantial improvements in both accuracy and speed. Across all scenarios and tasks, Copilot subjects experienced a 34.53% improvement in overall accuracy and a 29.79% reduction in task completion time. We also find that the productivity benefits vary by task type, with more complex tasks showing greater improvement. In free response tasks, Copilot users identified 146.07% more relevant facts and reduced task completion time by 61.14%. Subject satisfaction with Copilot was high, with participants reporting reduced effort and a strong preference for using the tool in future tasks. These findings suggest that GAI tools like Copilot can significantly enhance the productivity and efficiency of IT admins, especially in scenarios requiring information synthesis and complex decision-making.
@@ 1 Introduction
As generative AI (GAI) tools continue to mature, organizations are interested in understanding their impact on workforce productivity to guide adoption decisions. However, the effects of these tools can vary significantly depending on the specific domain and application. So, it is important to study each application individually. In this study, we evaluate the benefits of generative AI on information technology administrators (“IT admins”) by conducting a randomized controlled trial (RCT) with IT admins using Microsoft’s Security Copilot (“Copilot”). We find that Copilot confers substantial benefits to IT admins, including accuracy improvements across all scenarios and tasks (34.53%*** on average) and time savings across all scenarios (29.79%*** on average). [^1] Throughout, we indicate statistical significance using asterisks, where * indicates $0.05<p\leq 0.1$, ** indicates $0.01<p\leq 0.05$, and *** indicates $0<p\leq 0.01$.
@@
We recruited 182 subjects and asked them to sign into an environment created for this experiment. Half of the subjects were assigned to the control group and were granted access to the Entra and Intune admin centers. The other half were assigned to the treatment group and were additionally given access to Copilot within the Entra and Intune admin centers. Then, subjects used these tools and the simulated data in that environment to respond to our experimental tasks.
@@ 3 Methodology
We recruited our subject pool through Upwork, a marketplace for freelancers. We required subjects to be proficient in reading and writing English and to have a positive reputation on Upwork. We told them to expect the tasks to take approximately two hours to complete. We offered performance incentives for combined speed and accuracy. Specifically, we multiplied the percentiles of each subject’s speed and accuracy and awarded payments on the following scale:
@@ 4 Results
The results answer RQ1 in the affirmative: Copilot does offer speed and accuracy improvements in common IT scenarios. Across all tasks, we measure a 34.53%*** increase in accuracy and a 29.79%*** reduction in task completion time. In the subsections on accuracy and speed below, we present detailed results answering RQ1-RQ4.
@@ 4.1 Accuracy
Free Response Content 3 1.66 0.67 146.07%***
@@
Copilot shows statistically significant accuracy improvements in two out of three task types. The effect sizes generally track the task complexity, with the greatest benefits arising in free response tasks where Copilot subjects reported 146.07%*** more relevant facts than our control group. In contrast to MC and SATA tasks, which only require selecting the correct answers from a short list of possibilities, the free response tasks leverage AI’s unique ability to find and report relevant facts about a given scenario from a large and complex corpus, e.g., looking through sign-in logs to find and categorize sign-in failures caused by conditional access policies. Overall, Copilot yields an accuracy improvement of 21.76%*** on SATA tasks across the sign-in troubleshooting and device troubleshooting scenarios.
@@ 4.2 Speed
We measure statistically significant time savings in all scenarios. Table 3 summarizes these results by task type and scenario. The “Aggregate Results, All Content (Holding accuracy constant)” numbers represent our headline speed improvements, a 29.79%*** time savings for Copilot users. We also note that Copilot generally involves greater latency than typical page-loading in the admin centers, which is common among GAI tools. This necessarily slowed the Copilot users. Product improvements should reduce this latency and further increase the time savings for users with Copilot.
@@
All Content 53.84 44.29 63.10 -29.79%***
SATA 17.81 15.25 20.14 -24.28%***
MC 10.91 12.02 9.60 25.13%***
MC (Drop first) 7.17 7.47 6.71 10.92%
Free Response 30.98 17.03 43.82 -61.14%***
@@
All Content 24.04 16.63 30.46 -45.41%***
SATA 6.41 5.04 7.63 -33.98%***
Free Response Content 17.37 11.59 22.40 -48.23%***
@@
All Content 20.00 17.45 22.23 -21.64%***
MC 10.91 12.02 9.60 25.13%***
MC (Drop first) 7.17 7.47 6.71 10.92%
Free Response Content 12.17 5.43 19.27 -71.80%***
@@
We measure time savings for Copilot across all three scenarios. The time savings included 45.41%*** for sign-in troubleshooting, 21.64%*** for device policy management, and 15.69%*** for device troubleshooting.
@@
We find that Copilot’s time savings track its accuracy improvements across task types. That is, the biggest gains again come from free response (61.14%*** time savings), followed by SATA (24.28%***). Lastly, the MC questions show a decrease in speed for Copilot users (25.13%*** time increase). For these MC questions we also estimate a negative relationship between accuracy and time spent for the control group. This is likely because control subjects who knew the answer responded quickly, and those who did not know it gained little accuracy by spending more time by, for example, searching the internet. This can explain the data we see, where even though Copilot provided an advantage in responding accurately, controlling for accuracy in calculating its time savings does not make up for the disadvantage of its latency.
We also notice that Copilot users spent much more time than the control users on the first MC question compared to the subsequent three questions (see figure 2). We think this suggests Copilot users were learning how to use Copilot for this type of task. Therefore, we drop the first question and recompute the time differences across the remaining tasks, which results in no statistically significant difference in task completion time (see the “Drop first” entry in table 3).
@@
As we did with accuracy, we examine time savings across different experience levels and domains. Our findings once again reveal statistically significant point estimates of time savings for all groups, as shown in figure 3. However, we do not observe statistically significant differences in estimated effect sizes between any of the groups. The left panel illustrates Copilot’s impact on task completion times for the sign-in troubleshooting task, with a 50.1%*** and 45.2%*** time savings for Copilot users with 0-3 and 3+ years of experience, respectively. These point estimates are consistent with intuition and our accuracy findings.
The right panel of figure 3 shows time savings of 12.6%*** and 30.2%*** for 0-3 and 3+ years of experience, respectively, for device management tasks. This suggests that experienced network management subjects benefit more from Copilot in terms of time savings than their less experienced counterparts, which runs counter to intuition and expectation. This is driven by the fact that less experienced control subjects are driving the negative relationship between task completion time and accuracy for the device policy management MC question, mentioned above. This shrinks the possible time savings for the less experienced Copilot subjects.